# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v1

@@ Abstract
We present a fine-grained theoretical analysis of the performance gap between reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO) under a representation gap. Our study decomposes this gap into two sources: an explicit representation gap under exact optimization and an implicit representation gap under finite samples. In the exact optimization setting, we characterize how the relative capacities of the reward and policy model classes influence the final policy qualities. We show that RLHF, DPO, or online DPO can outperform one another depending on the type of model mis-specifications. Notably, online DPO can outperform both RLHF and standard DPO when the reward and policy model classes are isomorphic and both mis-specified. In the approximate optimization setting, we provide a concrete construction where the ground-truth reward is implicitly sparse and show that RLHF requires significantly fewer samples than DPO to recover an effective reward model—highlighting a statistical advantage of two-stage learning. Together, these results provide a comprehensive understanding of the performance gap between RLHF and DPO under various settings, and offer practical insights into when each method is preferred.
[…]
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF, Christiano et al. (2017); Ziegler et al. (2019)) is an important paradigm improving the natural language understanding and generation capabilities of large language models (LLMs). The core idea of RLHF is to utilize pair-wise comparison between responses from human annotators, as directly collecting absolute reward signals is hard. There are two stages in RLHF: the reward modeling stage and the policy optimization stage. The reward modeling stage assumes human preferences follow the Bradley-Terry (BT) model (Bradley and Terry, 1952), allowing a prompt-response pair to be assigned a scalar reward. Thus, a reward model $r_{\phi}$ could be trained using negative log-likelihood loss function from human preferences. In the policy optimization stage, the base LM is “online” fine-tuned with RL algorithms such as proximal policy optimization (PPO, Schulman et al. (2017)), based on $r_{\phi}$ under a Kullback-Leibler (KL) divergence-regularized multi-armed bandit setting. And the key assumption behind this two-stage pipeline is the realizability of the ground-truth reward.
@@
However, the assumptions of tabular parameterization and realizability often do not hold in practice, particularly when the reward model is significantly smaller than the policy model (e.g., $6$B vs. $175$B in Ouyang et al. (2022), indicating a clear disparity in representational capacity), when the policy model class is heavily mis-specified due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data. Consequently, one should not expect DPO to perform identically to RLHF. This motivates the central question of our investigation:
@@
To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($y$ is a response and we omit prompts): $V_{r^{\star}}^{\pi}:=\mathbb{E}_{y\sim\pi}[r^{\star}(y)]-\beta\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)$, where $\pi$ is a distribution over responses, $\pi_{{\mathsf{ref}}}$ is the base model. Let $\pi^{\star}:=\operatorname{argmax}_{\pi}V_{r^{\star}}^{\pi}$ be the ideal optimal policy.
@@
Sec. 3.1 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi)\propto-V_{r^{\star}}^{\pi}+\mathcal{O}(\delta^{2})$ Sec. 3.3 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ Sec. 3.2 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$ Sec. 3.4 It depends on
@@
$\bullet$ For approximate optimization, i.e., the finite-sample regime, we study the implicit representation gap incurred by statistical efficiencies in Section 4. We construct a simple task where the ground-truth reward to is a dual-token linear function with feature dimension $d$ and implicit sparsity $k$, and the total number of samples is $n$. Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, decreasing the error to $\mathcal{O}(\sqrt{k\log d/n})$. This result indicates that DPO is less sample-efficient than RLHF, leading to inferior performance.
Finally, we use empirical experiments to corroborate these theoretical findings in Section 5.
@@ 2 Preliminaries
[…]
Multi-armed Bandits and Policies. A multi-armed bandit is defined by an action space $\mathcal{Y}$ and a reward function $r:\mathcal{Y}\to\mathbb{R}$. A policy $\pi\in\Delta(\mathcal{Y})$ represents a probability distribution over the action space. Given the autoregressive mechanism of LM, it is natural to consider the prompt and response together and view the combination as a whole action. In this study, our results are presented in the context of multi-armed bandits.
Model class and value function. Let =$\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d_{R}}\}$ denote the reward model class, and $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{d_{P}}\}$ denote the policy model class, where $d_{R},d_{P}\in\mathbb{N}$. For a reward function $r$ and policy $\pi$, we define the regularized value function as:
$V_{r}^{\pi}:=\underset{y\sim\pi}{\mathbb{E}}[r(y)]-\beta\,\mathsf{KL}\left(\pi\|\pi_{{\mathsf{ref}}}\right)\;,$
where $\beta>0$ is the regularization coefficient, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy. Let $r^{\star}$ denote the ground-truth reward function, and $\pi^{\star}$ denote the optimal policy for $V_{r^{\star}}^{\pi}$. A well-known fact (Rafailov et al., 2023) is that $\pi^{\star}(y)=\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)/Z$, where $Z:=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y)\exp(r^{\star}(y)/\beta)$ is the partition function. The goal of preference-based policy learning is to find a policy $\pi_{\theta}\in\Pi$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$. We define the oracle value as $V^{\star}_{\Pi}:=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}$.
Bradley-Terry (BT) model. Given an implicit reward oracle $r:\mathcal{Y}\to\mathbb{R}$, Bradley and Terry (1952) assume that human preference distribution $p^{\star}:\mathcal{Y}\times\mathcal{Y}\rightarrow\Delta(\{0,1\})$ satisfies:
$p^{\star}(y_{1}\succ y_{2})=\sigma\left(r^{\star}(y_{1})-r^{\star}(y_{2})\right)\;.$
This means response $y_{1}$ is favored over $y_{2}$ with probability $p^{\star}(y_{1}\succ y_{2})\;$ by human annotators.
@@
$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma(r_{\phi}(y)-r_{\phi}(y^{\prime}))+p^{\star}(y^{\prime}\succ y)\log\sigma(r_{\phi}(y^{\prime})-r_{\phi}(y))\right]\;.$
In practice, people can have a pre-collected human preference dataset $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{N}$, where in the $i^{\text{th}}$ sample, $y_{w}^{(i)}\succ y_{l}^{(i)}$ means that $y_{w}^{(i)}$ is preferred over $y_{l}^{(i)}$. And the empirical loss is:
$\mathcal{L}_{\textup{RLHF}}(\phi)=-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\phi}(y_{w}^{(i)})-r_{\phi}(y_{l}^{(i)}))\;.$
Then using the reward model $r_{\textup{RLHF}}$, the policy learning stage returns $\pi_{\textup{RLHF}}=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\textrm{RLHF}}}^{\pi}$.
Direct approach of DPO. By leveraging the surrogate reward $\hat{r}_{\theta}(y):=\beta\log\frac{\pi_{\theta}(y)}{\pi_{{\mathsf{ref}}}(y)}$, DPO bypasses reward learning and directly learns the policy from preference data:
$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{y,y^{\prime}\sim\pi_{{\mathsf{ref}}}}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)+p^{\star}(y^{\prime}\succ y)\log\sigma\left(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y)\right)\right]\;.$
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}\right)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime})\log\sigma\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)+p^{\star}(y^{\prime}\succ y)\log\sigma\left(\hat{r}_{\theta}(y^{\prime})-\hat{r}_{\theta}(y)\right)\right]\;.$
3 Exact Optimization: Fine-grained Representation Gap Induced by Model Mis-specification
@@ 3.1 No Model Mis-specification
We begin with the fully realizable setting, where both the ground-truth reward function and the optimal policy lie within their respective model classes. Most prior works (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Swamy et al., 2025) focus on this setting.
@@
Both RLHF and DPO are capable of recovering the true optimal policy under ideal conditions. In this regime, RLHF directly optimizes $V_{r^{\star}}^{\pi_{\theta}}$ in the policy learning stage. Proof deferred to Section C.1.
@@
Shi et al. (2025) propose a sampling strategy to accelerate convergence in online DPO, and Feng et al. (2025) further refine this approach, showing its connection to the RLHF objective from a gradient-based perspective. Below, we show a result which is analogous to Theorem 4.1 in (Feng et al., 2025), but from the objective perspective rather than the gradient perspective.
@@
with a ratio $\alpha_{1}=1$ and $\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;$.
@@
To sample a pairwise data $(y_{1},y_{2})$ from PILAF sampler, we sample $y_{1}\sim\pi^{\mathsf{s1}},y_{2}\sim\pi^{\mathsf{s2}}$, w.p. $\alpha_{1}/(\alpha_{1}+\alpha_{2})$ from sampler ➀, and w.p. $\alpha_{2}/(\alpha_{1}+\alpha_{2})$ from sampler ➁.
@@
Given $R_{\max},\delta\in\mathbb{R}_{+}$, s.t. $0\leq r^{\star}(y)\leq R_{\max}$, $\forall y\in\mathcal{Y}$, and $|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|\leq\delta$, $\forall y,y^{\prime}\in\mathcal{Y}$, then with $\pi^{\mathsf{s}}$ defined in Definition 1, we have:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z(\theta)\right)}\left\{-V_{r^{\star}}^{\pi_{\theta}}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\right\}\;,$
where $\epsilon_{y,y^{\prime}}\in\mathbb{R}$ are noises s.t. $|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$ and $Z(\theta):=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))$ can be viewed as an adaptive step size.
@@
This result indicates that, with an appropriate sampler, the objective of online DPO can approximate the true value objective up to a second-order deviation. However, this deviation can become substantial when the reward function has a large scale or the ground-truth reward is poorly fitted. In such scenarios, the objective of online DPO may significantly deviate from the value function, leading to degraded convergence or even divergence. Proof deferred to Section C.2.

### v2

@@ Abstract
We present a fine-grained theoretical analysis of the performance gap between reinforcement learning from human feedback (RLHF) and direct preference optimization (DPO) under a representation gap. Our study decomposes this gap into two sources: an explicit representation gap under exact optimization and an implicit representation gap under finite samples. In the exact optimization setting, we characterize how the relative capacities of the reward and policy model classes influence the final policy qualities. We show that RLHF, DPO, or online DPO can outperform one another depending on type of model mis-specifications. Notably, online DPO can outperform both RLHF and standard DPO when the reward and policy model classes are isomorphic and both mis-specified. In the approximate optimization setting, we provide a concrete construction where the ground-truth reward is implicitly sparse and show that RLHF requires significantly fewer samples than DPO to recover an effective reward model—highlighting a statistical advantage of two-stage learning. Together, these results provide a comprehensive understanding of the performance gap between RLHF and DPO under various settings, and offer practical insights into when each method is preferred.
@@ 1 Introduction
Reinforcement learning from human feedback (RLHF, Christiano et al. (2017); Ziegler et al. (2019)) is an important paradigm improving the natural language understanding and generation capabilities of large language models (LLMs). The core idea of RLHF is to utilize pair-wise comparison between responses from human annotators, as directly collecting absolute reward signals is hard. There are two stages in RLHF: the reward modeling stage and the policy optimization stage. The reward modeling stage assumes human preferences follow the Bradley-Terry (BT) model (Bradley and Terry, 1952), allowing a prompt-response pair to be assigned a scalar reward. Thus, a reward model $r_{\phi}$ could be trained using negative log-likelihood loss function from human preferences. In the policy optimization stage, the base LM is “online” fine-tuned with RL algorithms such as proximal policy optimization (PPO, Schulman et al. (2017)), based on $r_{\phi}$ under a Kullback-Leibler (KL) divergence-regularized bandit setting. And the key assumption behind this two-stage pipeline is the realizability of the ground-truth reward.
@@
However, the assumptions of tabular parameterization and realizability often do not hold in practice, particularly when the reward model is significantly smaller than the policy model (e.g., 66B vs. 175175B in Ouyang et al. (2022), indicating a clear disparity in representational capacity), when the policy model class is heavily restricted due to limited computational resources, or when the reward model is sub-optimal owing to limited preference data. These situations are examples of model mis-specification, a common issue in practice due to limitations in model capacity or data. Consequently, one should not expect DPO to perform identically to RLHF under model mis-specifications. This motivates the central question of our investigation:
@@
To quantify the problem, we choose the performance metric as the expected value of the original regularized bandit problem using the ground-truth reward $r^{\star}$ ($x$ is a prompt, and $y$ is a response): $V_{r^{\star}}^{\pi}:=\mathbb{E}_{x\sim\rho}\left[\mathbb{E}_{y\sim\pi(\cdot|x)}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$, where $\rho$ is a pre-fixed distribution over prompts, $\pi$ is a distribution over responses, and $\pi_{{\mathsf{ref}}}$ is the base model. Let $\pi^{\star}:=\operatorname{argmax}_{\pi}V_{r^{\star}}^{\pi}$ be the ideal optimal policy.
@@
Sec. 3.1 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}=V_{r^{\star}}^{\pi_{\textup{DPO}}}=V_{r^{\star}}^{\pi^{\star}}$ $\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi(\cdot|x))\propto-V_{r^{\star}}^{\pi(\cdot|x)}+\mathcal{O}(\delta^{2})$ Sec. 3.3 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\leq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ Sec. 3.2 $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $V_{r^{\star}}^{\pi_{\textup{RLHF}}}\geq V_{r^{\star}}^{\pi_{\textup{DPO}}^{\textup{online}}}$ Sec. 3.4 It depends on
@@
$\bullet$ For approximate optimization, i.e., the finite-sample regime, we study the implicit representation gap incurred by statistical efficiencies in Section 4. We construct a simple task where the ground-truth reward to is a dual-token linear function with feature dimension $d$ and implicit sparsity $k$, and the total number of samples is $n$. Even without mis-specifications, we can reveal a separation between RLHF and DPO under this setting: the estimation error of DPO is $\Omega(d/n)$, while reward learning in RLHF can effectively leverage sparsity, decreasing the error to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$. This result indicates that DPO is less data-efficient than RLHF, leading to inferior performance.
Finally, we conduct numerical experiments to corroborate these theoretical findings in Section 5.
@@ 2 Preliminaries
Notation. Let $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ be the sigmoid function, where $\sigma(x)=1/(1+\exp(-x))$. For any set $\mathcal{X}$, $\Delta(\mathcal{X})$ represents the set of probability distributions over $\mathcal{X}$. $\mathsf{sg}\left(\right)$ is the stopping-gradient operator, where $\nabla_{\theta}[\mathsf{sg}\left(f(\theta)\right)]=\boldsymbol{0}$. Let $e_{k}$ be a one-hot vector with $1$ on its $k\textsuperscript{th}$ entry and 0 on other entries. We use $f(\theta){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}g(\theta)$ to indicate $\nabla_{\theta}f(\theta)=\nabla_{\theta}g(\theta)$.
Bandits and Policies. A bandit is defined by a state space $\mathcal{X}$, an action space $\mathcal{Y}$, and a reward function $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$. A policy $\pi:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$ represents a probability distribution over actions given a state. Note that, we sometimes omit the prompt $x$ for simplicity, so that $\pi\in\Delta(\mathcal{Y})$.
Model class and value function. Let $\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d_{R}}\}$ denote the reward model class, and $\Pi=\{\pi_{\theta}:\theta\in\mathbb{R}^{d_{P}}\}$ denote the policy model class, where $d_{R},d_{P}\in\mathbb{N}$. For a reward function $r$ and policy $\pi$, we define the regularized value function as:
$V_{r}^{\pi}:=\underset{x\sim\rho}{\mathbb{E}}\left[\underset{y\sim\pi(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\,\mathsf{KL}\left(\pi(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]\;,$
where $\beta>0$ is the regularization coefficient, and $\pi_{{\mathsf{ref}}}$ is a fixed reference policy. Let $r^{\star}$ denote the ground-truth reward function, and $\pi^{\star}$ denote the optimal policy for $V_{r^{\star}}^{\pi}$. A well-known fact (Rafailov et al., 2023) is that $\pi^{\star}(y|x)=\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)/Z(x)$, where $Z(x):=\sum_{y\in\mathcal{Y}}\pi_{{\mathsf{ref}}}(y|x)\exp(r^{\star}(x,y)/\beta)$ is the partition function. The goal of preference-based policy learning is to find a policy $\pi_{\theta}\in\Pi$ that maximizes $V_{r^{\star}}^{\pi_{\theta}}$. We define the oracle value as $V^{\star}_{\Pi}:=\max_{\pi\in\Pi}V_{r^{\star}}^{\pi}$.
Bradley-Terry (BT) model. Given an implicit reward oracle $r:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, Bradley and Terry (1952) assume that human preference distribution $p^{\star}:\mathcal{X}\times\mathcal{Y}\times\mathcal{Y}\rightarrow\Delta(\{0,1\})$ satisfies:
$p^{\star}(y_{1}\succ y_{2}|x)=\sigma\left(r^{\star}(x,y_{1})-r^{\star}(x,y_{2})\right)\;.$
This means response $y_{1}$ is favored over $y_{2}$ with probability $p^{\star}(y_{1}\succ y_{2}|x)\;$ by human annotators.
@@
$r_{\textup{RLHF}}=\underset{r_{\phi}\in\mathcal{F}}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma(r_{\phi}(x,y)-r_{\phi}(x,y^{\prime}))+p^{\star}(y^{\prime}\succ y|x)\log\sigma(r_{\phi}(x,y^{\prime})-r_{\phi}(x,y))\right]\;.$
Direct approach of DPO. By leveraging the surrogate reward $\hat{r}_{\theta}(x,y):=\beta\log\frac{\pi_{\theta}(x,y)}{\pi_{{\mathsf{ref}}}(x,y)}$, DPO bypasses reward learning and directly learns the policy from preference data:
$\pi_{\text{DPO}}=\underset{\pi_{\theta}\in\Pi}{\operatorname{argmax}}\;\underset{x\sim\rho;y,y^{\prime}\sim\pi_{{\mathsf{ref}}}(\cdot|x)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x))=-\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi^{\mathsf{s}}(\cdot|x)\right)}{\mathbb{E}}\left[p^{\star}(y\succ y^{\prime}|x)\log\sigma\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)+p^{\star}(y^{\prime}\succ y|x)\log\sigma\left(\hat{r}_{\theta}(x,y^{\prime})-\hat{r}_{\theta}(x,y)\right)\right]\;.$
3 Exact Optimization: Fine-grained Performance Gap Induced by Model Mis-specification
@@ 3.1 No Model Mis-specification
We begin with the fully realizable setting, where both the ground-truth reward function and the optimal policy lie within their respective model classes. While this assumption is often unrealistic in practice, it serves as a clean baseline and has been the main focus of most prior theoretical analyses (Xiong et al., 2024; Shi et al., 2025; Feng et al., 2025; Swamy et al., 2025).
@@
Both RLHF and DPO are capable of recovering the true optimal policy under ideal conditions. In this regime, RLHF directly optimizes $V_{r^{\star}}^{\pi_{\theta}}$ in the policy learning stage. Proof deferred to Section B.1.
@@
Although RLHF and DPO share a same solution, they differ in optimization trajectories and convergence rates. Shi et al. (2025) propose a sampling strategy to accelerate convergence in online DPO, and Feng et al. (2025) further refine this approach, showing its connection to the RLHF objective from a gradient-based perspective. Below, we show a result which is analogous to Theorem 4.1 in (Feng et al., 2025), but from the objective perspective rather than the gradient perspective.
@@
with a ratio $\alpha_{1}=1$ and $\alpha_{2}=\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\exp(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))\;$.
@@
Given a prompt $x$, we first randomly choose a sampler pair: select sampler ➀ w.p. $\alpha_{1}/(\alpha_{1}+\alpha_{2})$ and sampler ➁ otherwise. Then sample $y_{1}\sim\pi^{\mathsf{s1}}(\cdot\mid x)$ and $y_{2}\sim\pi^{\mathsf{s2}}(\cdot\mid x)$.
@@
Given $R_{\max},\delta\in\mathbb{R}_{+},x\in\mathcal{X}$, s.t. $0\leq r^{\star}(x,y)\leq R_{\max}$, $\forall y\in\mathcal{Y}$, and $|(r^{\star}(x,y)-r^{\star}(x,y^{\prime}))-(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))|\leq\delta$, $y,y^{\prime}\in\mathcal{Y}$, then with $\pi^{\mathsf{s}}$ defined in Definition 1, we have:
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}(\cdot|x)){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}(x)\right)}$ $\Bigg\{-\left[\underset{y\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}[r(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}(\cdot|x)\|\pi_{{\mathsf{ref}}}(\cdot|x)\right)\right]$
$+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}(\cdot|x)\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(x,y)-r^{\star}(x,y^{\prime})\right)-\left(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime})\right)\right]^{2}\right]\Bigg\}\;,$
where $\epsilon_{y,y^{\prime}}\in\mathbb{R}$ are noises s.t. $|\epsilon_{y,y^{\prime}}|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime(}R_{\max}+\delta)}$ and $Z_{\theta}(x):=\underset{y,y^{\prime}\sim\pi_{\theta}(\cdot|x)}{\mathbb{E}}1/\sigma^{\prime}(\hat{r}_{\theta}(x,y)-\hat{r}_{\theta}(x,y^{\prime}))$ can be viewed as adaptive step sizes for different prompts.
@@
This result indicates that, with an appropriate sampler, the objective of online DPO can approximate the true value function in prompt level. However, the second-order deviation can become substantial when $R_{\max}$ is large, or the ground-truth reward is poorly fitted. In such scenarios, the objective of online DPO may significantly deviate from the value function, leading to degraded convergence or even divergence. Proof deferred to Section B.2.